A telescope consists of two convex lens of focal length 16 cm and 2cm. What is angular magnification of telescope for relaxed eye ?

Grade:

1 Answers

Arun
4 years ago
Dear Manish
 
Angular magnification M = fo / fe = 16/2 = 8
 
Hope it helps
 
Regards
Arun ()askIITians forum expert